OpenStudy (anonymous):

Find the number of possible positive and negative real zeros of: 1. f(x)=x^4-x^3+2x^2+x-5 2. f(x)=3x^3-4x^2-5x+1 3. f(x)=5x^6+4x^5+3x^4-2x^3+2x^2-x-1

OpenStudy (jdoe0001):

well let us use the Descates Rule of Signs f(x)= x^4 - x^3 + 2x^2 + x - 5 + - + + - yes yes no yes so the sign changed 3 times, that means 3 or (3-2) real positive roots --------------------------------------- now let's check for f( -x ) f( -x )= x^4 + x^3 + 2x^2 - x - 5 no no yes no so the sign changed only once for f( -x ) meaning that, there's only 1 negative real root

OpenStudy (jdoe0001):

so that equation is a 4th degree polynomial, thus 4 roots it either has 3 real positive ones an 1 negative or 1 positive one and 1 negative, leaving room for 2 complex ones, for a total of 4

OpenStudy (anonymous):

Would the second one be 2 pos, 1 neg then?

OpenStudy (anonymous):

And the last one 4 pos, 2 neg?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

the 2nd one... let's see f(x)= 3x^3 - 4x^2 - 5x + 1 yes no yes 2 OR (2-2) real positive ones f( -x ) = -3x^3 - 4x^2 + 5x + 1 no yes no 1 real negative only yes 2 real positive and 1 negative # is a 3rd degree so 3 roots OR (2-2), 0 real positive ones and 1 negative and 2 complex ones

OpenStudy (anonymous):

ok perfect! and the third? @jdoe0001

OpenStudy (jdoe0001):

how many signs change did you get for f(x) and f ( -x ) ?

OpenStudy (anonymous):

I got 4 positive and 2 negative as my answer. is that correct?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, keep in mind that we're just counting when the sign hops from one term to the next and changes sign so \(\bf f(x)=5x^6+4x^5+3x^4-2x^3+2x^2-x-1 \\ \quad \\ \quad \\ f(-x)=5x^6-4x^5+3x^4+2x^3+2x^2+x-1\)

OpenStudy (jdoe0001):

for f( x ) I got 3 for f( -x ) another 3 too so the real positive ones are 3 OR (3-2) and the real negative ones are 3 OR (3-2) thus is a 6 degree, thus 6 roots 3 positive and 3 negative OR 3 positive and 1 negative and 2 complex OR 1 positive and 3 negative and 2 complex

OpenStudy (anonymous):

Thanks! :)

OpenStudy (jdoe0001):

yw